OpenStudy (anonymous):

A triangle has 2 angles that measure 60°. What kind of triangle is it? A. obtuse triangle B. right triangle C. scalene triangle D. equiangular triangle

563blackghost (563blackghost):

Well a triangle has the total of 180 degrees right? so try subtracting 180 by the total of the two angles. 60(2)-180=?

OpenStudy (xapproachesinfinity):

this is an equilateral triangle all angles are 60 degrees

OpenStudy (mathstudent55):

180 - 2(60) = 180 - 120 = 60 The third angle also measures 60 degrees. You have a triangle with all congruent angles measuring 60 degrees. Now you need to pick the correct answer.